OpenStudy (anonymous):

If a circle C with radius 1 rolls along the outside of the circle x^2+y^2=16, a fixed point P on C traces out a curve called an epicycloid, with parametric equations x=5cos(t)-cos(5t), y=5sin(t)-sin(5t). Find the area it encloses.

OpenStudy (anonymous):

Have you learned about polar coordinates?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

We are doing line integrals and greens theorem now though. Is there some way I could solve it using Green's Theorem?

OpenStudy (anonymous):

Do you remember what greens theorem states?

OpenStudy (anonymous):

\[A=\int\limits_{C}^{}x \ dy=-\int\limits_{C}^{}y \ dx= \frac{ 1 }{ 2 }\int\limits_{C}^{}x \ dy \ - y\ dx\]

OpenStudy (anonymous):

Ah, that's easy then.

OpenStudy (anonymous):

\[ \oint_C xdy = \int_a^bx\frac{dy}{dt}dt \]Where \(a,b\) are the begining and ending of the parametrization

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I remember what to do from here

OpenStudy (anonymous):

Ok actually, I know how to do this just plug in and such so that I have: But that will take forever to solve. Is there an easier way to solve this?

OpenStudy (anonymous):

I don't think your xdy/dt is correct

OpenStudy (anonymous):

I just checked and mathematica says that the dy/dt part I had was right?

OpenStudy (anonymous):

\[ (5\cos(t)-\cos(5t))\cdot (5\cos(t)-5\cos(5t)) \]

OpenStudy (anonymous):

Yep that's what I have

OpenStudy (anonymous):

What's so hard about that integral?

OpenStudy (anonymous):

You can do either one, but the first one seems easiest

OpenStudy (anonymous):

I got it thanks!